Consider a process consisting of five resources that are operated eight hours per day. The process works on three different products, A, B, and C;

Resource

Number of Workers

Processing Time for A (Min./Unit)

Processing Time for B (Min./Unit)

Processing Time for C (Min./Unit)

1

2

5

5

5

2

2

3

4

5

3

1

15

0

0

4

1

0

3

3

5

2

6

6

6

Demand for the three different products is as follows: product A, 40 units per day; product B, 50 units per day; and product C, 60 units per day.

(a) What is the average processing time in resource 2 (in minutes per unit)?
(b) What is the capacity of resource 2 (in units per day)?
(c) What is the demand for resource 2 in hours per day?
(d) What is the available resource time at resource 2 in hours per day?
(e) What is the implied utilization of resource 2?
Answer #1

Total processing time for Unit A , B & C = 3+4+5 = 12 minutes

Average processing time for 1 unit of each = 12 / 3 = 4 minutes

Average processing time in Resource 2 with 2 workers = 4 / 2 minutes per unit = 2 minutes / unit

b) Capacity of Resource 2 = Total time available in a day / Average processing time = 8 * 60 / 2 = 240 units / day

c) Demand for Resource 2 in hours per day = Sum of (Demand * Processing time) = 40 * 3 +50*4+5*60 = 620 minutes

or 620 / 60 = 10.33 hours

d) Available resource time at resource 2 in hours per day = 2 workers * 8 hours = 16 hours

e) Implied utilization = Demand / Available time = 10.33 / 16 = 64.58%